OpenStudy (anonymous):

solve by the eliminatin method 7r-9s=-1 9r+7s=73 the solution is

OpenStudy (anonymous):

multiply first equation by -9 & second - by 7; add together... let me know if there is a problem

OpenStudy (anonymous):

could you show me i do not know how to work this

OpenStudy (anonymous):

multiply first equation by -9 & second - by 7: -63r+81s=9 63r+49s=511 add together... should "eliminate" s, right...?

OpenStudy (anonymous):

will eliminate r (not s!)

OpenStudy (anonymous):

what did you got for s?

OpenStudy (anonymous):

after adding 2 equations together you should be left with: 130s=520 s=520/130 s=....?

OpenStudy (anonymous):

Is th answer 4

OpenStudy (anonymous):

good! now find r, please to do that, put value of s in ANY original equations above. Try it, please

OpenStudy (anonymous):

confused now so that was not the end of the equation

OpenStudy (anonymous):

in original equations you have 2 variables. we found s, now - need to find r

OpenStudy (anonymous):

because s=4, then by using original: 7r-9s=-1 we will have: 7r-9*4=-1 7r-36=-1 can you solve it for r?

OpenStudy (anonymous):

29

OpenStudy (anonymous):

7r=36-1 7r=35 r=35/7 r=5

OpenStudy (anonymous):

so, you got answer: r=5 and s=4 Good Luck!

OpenStudy (anonymous):

so the solution is an ordered pair of (5,4) correct

OpenStudy (anonymous):

right

OpenStudy (anonymous):

thanks you inik you broke it down very simple for me to understand I have not had math in over 35years thank you for your support

OpenStudy (anonymous):

you are very welcome!! let me know if there are any questions - happy to help :)

OpenStudy (anonymous):

okay